Determine the concentration (in mol / L) of a solution formed from 6.34 g of iron (III) borate and 150 ml of water.


1
Expert's answer
2021-06-07T03:16:06-0400

C(in moles)=V(substance)/V(solution)

n(FeBO3)=m(FeBO3)/M(FeBO3)=6.34g/115g/mol=0.06mol

V(FeBO3)=1.23l

V(solution)=1.23l+0.15l=1.38l

C=1.23l/1.38l=0.9mol